OpenStudy (anonymous):

What does decartes rules of signs tell you about the real roots of the polynomial? -x^3+5x^2-5x-6=0

OpenStudy (anonymous):

-x^3+5x^2-5x=6 -x((x-5)x+5)-6=0

OpenStudy (tkhunny):

As it stands -+-- That's two sign changes. 2, or 0 positive Real roots Switching to f(-x) +++- That's one sign change. There MUST be one negative Real root.